OpenStudy (anonymous):

What is the center and radius of the circle with the given equation? (x-1)^2+(y+1)^2=4.

OpenStudy (anonymous):

so would it be center (-1,1) radius 4? :)

OpenStudy (jdoe0001):

let's rewrite yours a bit $$\bf (x-h)^2+(y-k)^2=r^2\\ (x-1)^2+(y+1)^2=4 \implies (x-(\color{blue}{+1}))^2+(y-(\color{blue}{-1}))^2=2^2 $$

OpenStudy (anonymous):

okay, so it would be 1,-1 with a radius of four?

OpenStudy (jdoe0001):

\(\bf r^2 = 4\\ r = \sqrt{4} = 2\)

OpenStudy (anonymous):

I didnt know we were supposed to square it back, thank you so much!

OpenStudy (jdoe0001):

yw